OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that is perpendicular to the given line and that passes through the given point. 3x-4y=8;(-6,5)

OpenStudy (jjuden):

3y = -4x-9, or y = -4x/3 - 3.

OpenStudy (texaschic101):

slope intercept form is y = mx + b first we need to find the slope of the given equation. Do you know how to do that ?

OpenStudy (texaschic101):

first we put the equation in y = mx + b form and the slope will be in the m position. 3x - 4y = 8 --- subtract 3x from both sides -4y = -3x + 8 -- now divide both sides by -4 y = 3/4x - 2....the slope of this equation is 3/4. However, we are looking for a perpendicular line...so we need the negative reciprocal slope. All that means is " flip " the slope and change the sign. So the slope we need is -4/3...see how I flipped the slope and changed the sign. Now we use y = mx + b slope(m) = -4/3 (-6,5)...x = -6 and y = 5 now we sub and since we already know slope(m), we need to find y intercept (b). 5 = -4/3(-6) + b 5 = 8 + b 5 - 8 = b -3 = b therefore, the perpendicular equation is : y = (-4/3)x - 3

OpenStudy (texaschic101):

if you have any questions, please ask :)

OpenStudy (anonymous):

Thank you!